The Kyu Lee provides you with the following miscellaneous data regarding operations in 20X0 (in thousands of South Korean won, ₩):

Break-even point in sales

₩ 40,000

Direct material used

  29,000

Gross Profit

  20,000

Contribution Margin

    25,000

Direct Labour

    30,000

Sales

100,000

Variable manufacturing overhead

      5,000

There are no beginning or ending inventories

Compute (a) the fixed manufacturing overhead (b) variable selling and administrative expenses and (c) fixed selling and administrative expenses
Answer #1

a) Cost of goods sold = Sales-Gross profit = 100000-20000 = 80000

Fixed manufacturing overhead = 80000-29000-30000-5000 = 16000

b) Variable cost = Sales-Contribution margin = 100000-25000 = 75000

Variable selling and administrative expense = 75000-29000-30000-5000 = 11000

c) Contribution margin ratio = 25000/100000 = 25%

Fixed cost = 40000*25% = 10000

Fixed selling and administrative overhead =